Please help me solve this equation:
Find the slope and the y-intercept of 5y+6x=5. Then graph the equation.

Answer by nadinadan(32)

5y+6x=5 To solve this equation we need to write this in the correct form which is y=mx+b
so 5y=-6x+5 so y= -6x/5 + 1
the slope is m and the y-intercept is b
so in our case m=-6/5 and y-intercept is 1 for our equation.